### v1

# Depth to Anatomy: Learning Internal Organ Locations from Surface Depth Images

Abstract
Automated patient positioning plays an important role in optimizing scanning procedure and improving patient throughput. Leveraging depth information captured by RGB-D cameras presents a promising approach for estimating internal organ positions, thereby enabling more accurate and efficient positioning. In this work, we propose a learning-based framework that directly predicts the 3D locations and shapes of multiple internal organs from single 2D depth images of the body surface. Utilizing a large-scale dataset of full-body MRI scans, we synthesize depth images paired with corresponding anatomical segmentations to train a unified convolutional neural network architecture. Our method accurately localizes a diverse set of anatomical structures, including bones and soft tissues, without requiring explicit surface reconstruction. Experimental results demonstrate the potential of integrating depth sensors into radiology workflows to streamline scanning procedures and enhance patient experience through automated patient positioning. The implementation and pretrained models are publicly available at https://github.com/EytanKats/orgloc.
keywords:
Radiology workflow, Organs localization
doi:
10.59275/j.melba.2024-AAAA
volume: 2
1 Introduction
[…]
The growing demand for radiology examinations underscores the need to optimize scanning procedures to reduce patient wait times and improve operational efficiency. A critical, yet time-intensive, aspect of this process is the patient preparation and planning phase, which is essential for optimizing image quality and diagnostic accuracy (van2020cinderellas). This phase typically involves manual positioning of the patient followed by scout imaging to facilitate detailed geometric planning for subsequent diagnostic scans. In standard clinical practice, a radiologic technologist initially aligns the patient on the table, often relying on external anatomical landmarks or laser positioning guides to approximate the region of interest. Once the patient is positioned inside the bore, a set of low-resolution, multi-planar scout images is acquired. These images are then used to visually confirm the patient’s position and to manually plan the geometry of the diagnostic sequences, such as slice orientation, field-of-view, and scan range (koken2009towards).
Despite recent advances in radiology hardware and software, the automation of patient positioning remains a largely underexplored area compared to other aspects of scan acceleration or reconstruction (shafieizargar2023systematic; heckel2024deep). Traditional workflows still rely heavily on the experience and judgment of technologists. This manual planning process is both operator-dependent and time-consuming, particularly in complex examinations involving multiple regions or when precise anatomical targeting is required. Furthermore, variability in technologist expertise and patient anatomy can lead to inconsistent positioning, necessitating repeat localizers or even additional diagnostic scans, thereby increasing overall examination time and reducing scanner throughput (danilouchkine2005operator).
Figure 1: Overview of the proposed automated patient positioning workflow. A depth sensor installed above the scanner bore captures a depth image of the patient lying on the table. This image is processed by a trained model that predicts the 3D bounding box of the target organ. The scanner table is then automatically adjusted to align the organ of interest with the center of the imaging field-of-view, reducing the need for manual repositioning and scout scans.
Automated patient positioning systems, capable of accurately identifying and localizing the anatomical region of interest, hold considerable promise for streamlining radiology workflows by automatically adjusting the patient table to the optimal position. By reducing reliance on manual table adjustments and anatomical landmark estimation, these systems can minimize inter-operator variability, decrease setup time, and mitigate the risk of positioning errors, improving both imaging efficiency and diagnostic consistency across radiology departments.
Among the emerging technological approaches in this area, RGB-D camera-based systems have shown notable potential in enhancing patient setup and positioning (booij2021automated; booij2019accuracy). These integrated systems use depth-sensing to generate surface maps of the patient’s body. By analyzing these surface contours, it becomes possible to detect and localize key anatomical landmarks, which serve as proxies for internal anatomical regions of interest. This information can then be used to infer the optimal imaging volume and adjust the patient table position automatically prior to scan acquisition, without the need for scout imaging. For example, incetan2020rgb demonstrated the utility of RGB-D sensors in automatically localizing target regions such as the head and thorax with high spatial accuracy. Similarly, karanam2020towards employed depth-aware landmark detection models to support robust, contactless patient positioning in MRI and CT workflows.
In the proposed workflow, a depth sensor is integrated into the room above the scanner bore to capture a depth image of the patient lying on the table. This depth image is then processed by the trained model, which outputs the 3D coordinates of the bounding box enclosing the organ of interest. Based on this prediction, the scanner table can be automatically adjusted to ensure that the relevant anatomy is optimally centered within the imaging field-of-view, minimizing the need for manual repositioning or iterative scout imaging (Figure 1). This approach offers a contactless, fast, and standardized method of patient localization, which could be particularly beneficial in high-throughput clinical environments or for procedures requiring precise anatomical targeting.
To ensure robust performance across a wide range of anatomical variations and body types, we train our model on a large-scale dataset consisting of approximately 10,000 whole-body MRI scans from the German National Cohort (NAKO) study (bamberg2022whole). This dataset includes subjects of diverse ages, body sizes, and physiological characteristics, enabling the model to generalize effectively across different patient populations. Our experimental evaluation demonstrates the effectiveness of the proposed method in accurately localizing 41 distinct internal anatomical structures, including bones and soft tissue organs, using only depth images.
The key contributions of this work are as follows:
  1. 1.
    We demonstrate the feasibility and effectiveness of a learning-based framework for estimating 3D locations of internal anatomical structures solely from depth images of the patient’s body surface.
  2. 2.
    We introduce a convolutional neural network architecture, Pixel-to-Voxel (Pix2Vox), which maps 2D depth images to 3D volumetric segmentation masks, enabling organ localization from a single surface view.
  3. 3.
    We validate the generalizability of our model across a wide range of body types and anatomical variations by training and evaluating on a large-scale, population-level dataset comprising approximately 10,000 whole-body MRI scans.
2 Related Works
The prediction of internal anatomical structures from external body surface representations has gained increasing attention due to its potential to enable non-invasive, efficient, and patient-specific modeling. Early work by wu2018towards introduced a learning-based framework that generates volumetric anatomical phantoms of skeletal and pulmonary structures from triangulated meshes of the patient’s external surface. This study demonstrated the feasibility of inferring detailed 3D internal anatomy from external geometry using deep models trained on co-registered body surface and internal organ data.
Intermediate 3D representations for internal anatomy estimation.
Several approaches address the prediction of internal anatomy by first deriving an intermediate 3D representation from depth images, which is then used to infer internal structures. Building on the integration of statistical human body models with internal anatomy, recent research has extended the widely adopted Skinned Multi-Person Linear (SMPL) model (loper2015smpl) to incorporate internal organ representations. The original SMPL is a parametric, skinned 3D body model that encodes human pose and shape with a low-dimensional set of learned parameters, enabling the synthesis of realistic body meshes through linear blend skinning. Guo et al. (guo2022smpl) proposed SMPL-A, which augments SMPL with deformable internal organ geometries, allowing organ shapes to adapt dynamically with body pose. While effective in modeling pose-aware internal anatomy, SMPL-A requires a patient-specific full-body mesh to initialize internal organ shapes, limiting its generalizability and real-time applicability in clinical environments. Similarly, OSSO (keller2022osso) and BOSS (shetty2023boss) infer internal organ shapes and locations by optimizing anatomical plausibility within the SMPL framework based on external body surface meshes. Extending this paradigm, the HIT (Human Internal implicit Tissues) framework (keller2024hit) employs a deep implicit function to predict continuous volumetric tissue distributions conditioned on a parametric body surface mesh. HIT leverages large-scale synthetic datasets and a canonical space aligned with SMPL to infer anatomically coherent internal structures. Despite promising results, these SMPL-based approaches depend on accurate surface mesh reconstruction and template fitting, which are computationally intensive. The LOOC framework (henrich2025looc) introduces an implicit volumetric model to infer internal organ locations and approximate shapes from depth-derived surface point clouds.
Prediction of 2D organ projections.
Recent studies by geissler2025predicting and kats2025internal successfully utilize convolutional neural networks (CNNs) that operate directly on depth maps to predict 2D projections of organs in the coronal plane. These works demonstrate that CNNs can effectively extract spatial cues from external body surfaces to infer accurate organ locations.
Prediction of 3D features from 2D depth images.
Teixeira et al. (teixeira2023automated) proposed a CNN-based method that directly estimates critical scan planning parameters—such as lung centroids and patient isocenter—from raw depth images. This approach bypasses SMPL-based registration and operates solely on depth data, providing a lightweight and robust solution for specific imaging tasks such as isocenter calculation in thoracic MRI.
In contrast to previous work on organ localization, our method introduces a lightweight, end-to-end framework that directly predicts explicit 3D bounding boxes and estimates shapes for 41 internal anatomical structures from a single depth image. Leveraging a unified Pix2Vox convolutional architecture, our approach eliminates the need for intermediate steps such as 3D surface reconstruction or canonical space registration, enabling rapid inference and practical integration into clinical workflows. To foster reproducibility and facilitate further research, our code and pretrained models are publicly available at https://github.com/EytanKats/orgloc.
3 Materials and Methods
Figure 2: Overview of the proposed hybrid Pix2Vox network architecture. Left: The full model consists of a 2D encoder, 3D decoder, and custom 2D-to-3D conversion layers that connect the encoder and decoder at the bottleneck and each skip connection. Right: A detailed view of the 2D-to-3D conversion block. This module reshapes the 2D feature maps by unsqueezing a feature dimension and applying a 3D convolution to produce a volumetric representation suitable for further 3D processing. For clarity, the batch dimension is omitted from all illustrated tensor shapes.
In this study, we leverage 10,020 full-body MRI scans from the National Cohort Study of Germany (NAKO) dataset (bamberg2022whole). This large and diverse dataset enables us to capture a broad spectrum of anatomical variability across age, sex, and body habitus, providing a robust foundation for training generalizable models of internal anatomy. Since paired depth images and organ segmentations are not natively available in the dataset, we synthetically generate training pairs from the MRI data.
To create depth images, we simulate depth sensor output by rendering the outer contours of the patient’s body from MRI volumes, mimicking the top-down perspective of real-world RGB-D sensors typically mounted above the scanner table. Simultaneously, we apply state-of-the-art automated segmentation tools to extract ground-truth 3D labels for 41 internal anatomical structures, including bones and soft tissue organs (Section 3.1).
The resulting dataset, composed of 2D depth images paired with corresponding 3D anatomical segmentations, serves as the training input for our proposed learning pipeline. We devised CNN and train Pix2Vox convolutional neural network, which maps each 2D depth image to a 3D volumetric prediction of the target organ bounding boxes Section 3.2).
3.1 Data Preparation
To simulate depth images from MRI volumes, we begin by normalizing each 3D MRI scan to the intensity range $[0,1]$. Next, we apply a global intensity threshold of 0.02 to generate a binary mask that captures the patient’s body. Small noise components and isolated artifacts are subsequently removed using binary morphological opening with a spherical structuring element, producing a cleaner mask. We then extract 2D orthographic depth projections in the coronal plane, simulating the top-down view of a depth sensor positioned above the patient. Specifically, for each pixel in the image plane, the depth value is computed as the closest point of intersection along the anterior-posterior axis with the segmented patient’s body. The resulting depth maps are normalized again to the range $[0,1]$. To suppress non-anatomical structures such as the patient table and background noise, we apply a depth cutoff, setting all pixel values below 0.3 to 0. This effectively removes table artifacts and regions far below the simulated camera plane. Finally, to refine the depth image, we apply grayscale morphological opening, which smooths remaining artifacts while preserving significant anatomical features such as limb contours and body curvature. The resulting depth images offer a standardized approximation of the patient’s body surface from a fixed sensor viewpoint and serve as input to our learning framework.
To generate high-quality anatomical reference labels, we leveraged recent state-of-the-art automated whole-body MRI segmentation tools, including TotalSegmentatorMRI (d2024totalsegmentator), MRSegmentator (hantze2024mrsegmentator), and TotalVibeSegmentator (graf2024totalvibesegmentator). These models, trained on large-scale MRI datasets, are capable of accurately segmenting a wide range of anatomical structures with high spatial precision. To further increase the number of segmented labels, the MR volumes were converted into synthetic CT volumes to utilize segmentation models available for CT data (wasserthal2023totalsegmentator; jaus2023towards), following the approach described in mensing2025mr. An aggregation procedure was implemented to select, for each anatomical structure, the segmentation mask with the highest accuracy. When a structure was available from multiple models, MR-based segmentations were prioritized over CT-based ones. The MRI segmentation models were ranked according to the similarity of their training data to our dataset, in the following order: TotalVibeSegmentator, MRSegmentator, and TotalSegmentatorMRI. For structures exclusively available from CT-based models, TotalSegmentator was preferred over Atlas DAP. The resulting aggregated masks encompass 137 distinct structures, including organs, vessels, bones, and tissue types.
Given the varying quality of automated segmentations across different structures and the need for relevance to clinical workflow optimization, we performed a careful selection process. In consultation with domain experts, and through extensive visual inspection of the output masks, we selected 41 anatomical structures (Table 1). These included bones and soft-tissue organs that are both (1) of practical importance for guiding scan planning and patient positioning, and (2) reliably segmented across a broad range of patients. To further improve the consistency of the labels and reduce artifacts introduced by the segmentation models, we applied a series of post-processing steps. These included hole filling to restore anatomical completeness, and connected components analysis to eliminate spurious isolated regions or misclassified voxels. These refinements ensured cleaner and more robust ground-truth volumes for training the model.
Figure 3: Mean absolute Detection Offset Error (DOE) averaged across all organs for each bounding box side (left-right, anterior-posterior, superior-inferior). The CNN models demonstrates strong localization accuracy, particularly in the coronal (left-right) and vertical (superior-inferior) dimensions, significantly outperforming the mean model baseline. They achieve sub-10mm average error in the coronal plane. Performance degrades along the anterior-posterior axis due to limited depth cues available from the coronal input image. Standard deviation is depicted as error bars extending in the positive direction only.
Table 1: Selected organ labels grouped by tissue type
Soft Tissue Organs Bones
spleen scapula left vertebrae_T12
kidney right scapula right vertebrae_T11
kidney left clavicula left vertebrae_T10
liver clavicula right vertebrae_T9
stomach femur left vertebrae_T8
pancreas femur right vertebrae_T7
lung right hip left vertebrae_T6
lung left hip right vertebrae_T5
trachea sacrum vertebrae_T4
thyroid gland vertebrae L5 vertebrae_T3
duodenum vertebrae L4 vertebrae_T2
urinary bladder vertebrae L3 vertebrae_T1
aorta vertebrae L2
heart vertebrae L1
3.2 Pix2Vox Model
For training, we employ a U-shaped convolutional neural network architecture that takes as input 2D depth images representing the patient’s body surface and outputs 3D volumetric segmentation masks of internal anatomical structures (Figure 2). The network is designed to bridge the dimensional gap between 2D input and 3D output using a combination of standard 2D encoder and 3D decoder modules, connected via custom 2D-to-3D conversion layers placed at the bottleneck and each skip connection.
The encoder consists of a series of convolutional and pooling layers that extract increasingly abstract spatial features from the input depth image. These features are encoded as 2D feature maps of shape $b\times c\times h\times w$, where $b$ denotes the batch size, $c$ the number of channels, and $h$, $w$ the spatial height and width, respectively.
To transition from 2D to 3D, we introduce conversion layers that reshape and transform the encoded 2D features into a format suitable for the 3D decoder. Specifically, for a given layer $l$, we apply the following transformation:
  1. 1.
    First, we unsqueeze the 2D feature maps along a new feature axis, producing a 5D tensor of shape $b\times 1\times c\times h\times w$. This tensor is interpreted as a volumetric feature map with a singleton depth dimension and channel depth equal to $c$.
  2. 2.
    Next, we apply a 3D convolutional layer with kernel size $k_{l}\times 1\times 1$, stride $s_{l}\times 1\times 1$, and output feature size $f_{l}$. The $k_{l}$, $s_{l}$, and $f_{l}$ parameters are chosen individually for each layer $l$ to produce an appropriate depth dimension and number of features for input into the 3D decoder ( Figure 2).
The decoder is composed of standard 3D convolutional layers with upsampling operations to progressively reconstruct the full volumetric segmentation output. The decoder also incorporates skip connections from the encoder, each of which passes through its corresponding 2D-to-3D conversion layer to ensure dimensional compatibility.
This hybrid 2D-3D design enables efficient processing of the 2D input depth images while producing dense 3D predictions, balancing memory efficiency with spatial expressiveness.
4 Experiments and Results
Figure 4: Dice coefficient for volumetric organs segmentation. Values are averaged for bilateral organs (e.g., kidneys, lungs) and vertebral subgroups (thoracic and lumbar vertebrae). Pix2Vox significantly outperforms the baseline, demonstrating its ability to generate accurate, patient-specific organ segmentations from a single coronal depth image.
To train our model, we randomly selected 8,528 pairs of depth images and their corresponding segmentation masks from the preprocessed dataset ( Section 3.1). A separate validation set of 494 samples was used to monitor training performance and guide model selection and hyperparameter tuning. For evaluation, we employed a dedicated test set of 998 samples, which was held out during training. This large and diverse set of examples enables a thorough assessment of the model’s performance and provides strong evidence of its ability to learn robust and generalizable features across range of anatomical variations.
4.1 Training and evaluation details
We train our model in a multi-label segmentation setting, applying a sigmoid activation function at the final layer to independently predict the presence of each anatomical structure. The loss function is a weighted sum of Dice loss and binary cross-entropy loss, with both components equally weighted at 0.5. To enhance the model’s robustness to anatomical variability and patient positioning, we apply data augmentations including random shifts, scaling, and rotations during training. This encourages the model to generalize across a broader range of spatial configurations. Training is performed using the Adam optimizer with an initial learning rate of 0.001. We apply a linear warm-up strategy over the first 1,000 training steps to stabilize early training, followed by cosine annealing to gradually reduce the learning rate to zero over the remaining steps. We use a batch size of 8, and training is carried out for 15,000 steps in total. To further improve robustness and reduce prediction noise during evaluation, we apply a post-processing step based on connected component analysis. This removes small, isolated predictions that are spatially distant from the main predicted regions, thereby reducing the likelihood of large errors caused by spurious or misclassified voxels.
Figure 5: Average Symmetric Surface Distance (ASSD) for volumetric organs segmentation. Values are averaged for bilateral organs (e.g., kidneys, lungs) and vertebral subgroups (thoracic and lumbar vertebrae). Pix2Vox exhibits markedly lower surface distances compared to the mean model, indicating superior boundary accuracy and shape fidelity.
4.2 Baselines
Mean model.
Given the well-aligned patient positioning in the NAKO MRI dataset and the consistent anatomical layout across subjects, we exploit this property to construct a mean segmentation baseline model. Despite variability in patient-specific characteristics such as age, height, and body shape, the high degree of spatial alignment in the dataset enables the generation of an average anatomical representation from the training data. To build this baseline, we aggregate all ground truth segmentations from the training set by computing the voxel-wise mean across all examples. Specifically, each label channel is summed across all training subjects, divided by the total number of samples, and then thresholded at 50% of the maximum voxel intensity for that label to obtain a binary segmentation. This process yields a static, patient-agnostic model that serves as a reference for evaluating the effectiveness of learning-based methods. Although simplistic, this mean model provides a strong benchmark in settings with highly standardized acquisition protocols, offering insights into how much a learning-based model improves upon pure spatial prior knowledge.
2.5D CNN.
Inspired by prior work by geissler2025predicting and kats2025internal, which leverage 2D CNNs to predict coronal plane projections of anatomical structures, we extend this concept to a more comprehensive 2.5D setting. Specifically, we employ separate 2D CNN models to predict the organ projections onto all three principal anatomical planes: coronal, sagittal, and axial. Each model is trained independently to generate a binary mask representing the projection of target anatomical structures in its respective plane, using the same depth image as input. By aggregating the outputs of these three models, we can infer 3D bounding boxes for each organ by combining the spatial extents across all dimensions. This approach offers a computationally efficient alternative to full 3D segmentation, while still capturing spatial context through multi-planar analysis.
Figure 6: Qualitative comparison between predicted and ground truth volumetric segmentations. The Pix2Vox model demonstrates a strong ability to accurately infer both the shape and location of internal organs solely from a single depth image. The examples highlight successful predictions for both bony (vertebrae and scapula) and soft tissue organs (lung and heart) across different organ sizes and positions.
4.3 Metrics
We evaluate the localization and segmentation performance of all models using a combination of complementary metrics that capture both spatial accuracy and anatomical shape fidelity. Bounding boxes are derived from the volumetric segmentation masks produced by the mean model, the Pix2Vox model, and the three-plane predictions of the 2.5D CNN. To quantify localization accuracy, we calculate the Mean Absolute Detection Offset Error (DOE), defined as the absolute distance between the corresponding faces of the predicted and ground truth bounding boxes along each axis. To better understand model behavior in different spatial dimensions, DOE is reported separately for each side of the bounding box (e.g., left, right, anterior, posterior, superior, inferior). This directional breakdown allows us to assess how accurately each method positions the anatomical structures along the full 3D extent. Importantly, this metric reflects practical constraints in clinical imaging scenarios, particularly in automated patient positioning and scan planning, where under- or over-estimating organ boundaries can lead to suboptimal coverage or increased scan time. To assess the model’s ability to reconstruct organ shape, we compute the Average Symmetric Surface Distance (ASSD) between the predicted and ground truth segmentations, which captures discrepancies at the organ boundary. In addition, we report the Dice coefficient as a standard volumetric overlap metric. Both Dice and ASSD are calculated only for models that produce full 3D volumetric predictions, enabling us to evaluate how well the organ shape can be recovered from 2D depth image inputs alone.
4.4 Results
Quantitative analysis using the Detection Offset Error (DOE) highlights the strong performance of the proposed Pix2Vox model in localizing internal organs in 3D space based solely on a single coronal depth image (Figure 3). Specifically, in the coronal plane, Pix2Vox demonstrates a more than twofold improvement in mean absolute DOE compared to the mean model baseline, achieving an average localization error of less than 10mm across all evaluated organs. However, performance degrades when localizing organs along the anterior-posterior (depth) axis, where the model shows higher variance and larger prediction errors. This limitation is not unexpected, given that the coronal depth image inherently provides less depth-specific information. Nonetheless, the ability of the CNN to infer organ depth from contextual and structural cues in a single 2D image remains noteworthy, demonstrating the model’s strong capacity for generalization. The 2.5D CNN approach, which integrates predictions from coronal, sagittal, and axial projections, exhibits superior localization performance in the coronal plane and achieves more robust bounding box estimation in lateral and vertical dimensions. However, it still underperforms in predicting organ position along the depth axis. This shortfall may be attributed to the limited representational capacity of the 2D projection-based architecture and the absence of a fully 3D decoding mechanism, which constrains the model’s ability to jointly reason across spatial dimensions.
Quantitative evaluation using the Dice coefficient and ASSD)(Figure 4 and Figure 5), alongside qualitative visualizations of Pix2Vox predictions (Figure 6), reveals the model’s remarkable ability to infer internal organ shape from a single 2D depth image. Unlike the mean model baseline, which simply reflects anatomical averages, Pix2Vox demonstrates a substantial improvement in both overlap and surface accuracy metrics. This performance gap highlights the model’s capacity to move beyond population-level anatomical assumptions and instead learn spatially-aware, patient-specific volumetric representations directly from depth cues present in surface imagery. The model’s strong performance across diverse organ types, sizes, and locations further supports the idea that it has learned robust spatial priors, enabling generalization across a heterogeneous population.
These findings suggest that internal organ anatomy can be successfully estimated directly from body surface depth information, highlighting its value as a non-invasive surrogate for internal imaging. This demonstrates the potential of integrating camera-based organ localization into radiological workflows, where such technology could enhance scan planning and optimize patient positioning.
5 Discussion
In this study, we introduce an approach for automated patient positioning that leverages body surface depth images to estimate the 3D position and shape of internal organs. Our quantitative and qualitative results demonstrate the feasibility of using a single depth image to approximate the internal anatomical layout. The proposed Pix2Vox model successfully learns to extract cues from surface morphology, such as torso shape, curvature, and contour depth, to predict the volumetric extent and spatial orientation of internal organs.
The model’s strong localization accuracy across multiple organs a promising direction to support scan planning and optimize table positioning without the need for external markers or scout imaging. Automated positioning could reduce reliance on trial-and-error adjustments by radiology technicians, potentially decreasing scan time, improving reproducibility, and reducing inter-operator variability. Additionally, the model’s capacity to generalize across diverse patient anatomies, as evidenced by its performance on a large and heterogeneous test set, underlines its robustness in dealing with physiological variation.
Despite these encouraging findings, several limitations and opportunities for future research must be acknowledged:
  • Domain gap between synthetic and real-world depth data: The depth images used in this study were synthetically generated from MRI-derived segmentations and surface reconstructions, which may not fully capture the variability present in real-world depth camera outputs. In clinical environments, additional sources of noise, such as occlusions, limited spatial resolution, lighting artifacts, and sensor-specific distortions, can affect image quality. Furthermore, patients may be covered with blankets, wear hospital gowns, or use assistive devices, all of which can obscure key body surface features. Bridging this domain gap is crucial for real-world applicability and may require domain adaptation techniques or model fine-tuning using real sensor-acquired data.
  • Assumption of consistent patient pose and orientation: The training data used in this work consisted of patients in a standardized supine position with minimal variation in posture. Although position-based data augmentations were employed during training to increase robustness, real-world scenarios often involve a wider range of patient poses due to comfort needs or mobility restrictions. The current model may be sensitive to these variations, and further research is needed to evaluate and improve its generalizability to diverse body orientations and clinical settings.
[…]
    We gratefully acknowledge the financial support by German Research Foundation: DFG, HE 7364/10-1, project number 500498869.
[…]
    The work follows appropriate ethical standards in conducting research and writing the manuscript, following all applicable laws and regulations regarding treatment of animals or human subjects.
[…]
    We declare we do not have conflicts of interest.
[…]
    Data from the German National Cohort (NAKO) are available for research purposes upon registration and formal request through the NAKO Research Platform: https://nako.de/en/research/. Access is subject to approval by the NAKO Data Use and Access Committee and compliance with data protection regulations.

### v2

@@
-Depth to Anatomy: Learning Internal Organ Locations from Surface Depth Images
+
 Abstract
-Automated patient positioning plays an important role in optimizing scanning procedure and improving patient throughput.
-Leveraging depth information captured by RGB-D cameras presents a promising approach for estimating internal organ positions, thereby enabling more accurate and efficient positioning.
-In this work, we propose a learning-based framework that directly predicts the 3D locations and shapes of multiple internal organs from single 2D depth images of the body surface.
-Utilizing a large-scale dataset of full-body MRI scans, we synthesize depth images paired with corresponding anatomical segmentations to train a unified convolutional neural network architecture.
-Our method accurately localizes a diverse set of anatomical structures, including bones and soft tissues, without requiring explicit surface reconstruction.
-Experimental results demonstrate the potential of integrating depth sensors into radiology workflows to streamline scanning procedures and enhance patient experience through automated patient positioning.
+Automated patient positioning can improve radiology workflow efficiency by reducing the time required for manual table adjustments and scout-based scan planning.
+We propose a learning-based framework that predicts 3D organ locations and shapes for 41 anatomical structures, including both bones and soft tissues, directly from a single 2D depth image of the body surface.
+Leveraging $10,020$ whole-body MRI scans from the German National Cohort (NAKO) dataset, we synthetically generate depth images paired with anatomical segmentations to train a convolutional neural network for volumetric organ prediction.
+Our method achieves a mean dice similarity coefficient of $0.44\pm 0.2$ and and a symmetric average surface distance of $7.69\pm 5.68$ mm across all structures.
+Furthermore, the model derives organ bounding boxes with a mean absolute detection offset of $10.99\pm 5.54$ mm.
+Qualitative results on real-world depth images confirm the ability of the model to generalize to practical clinical settings.
+These findings suggest that depth-only organ localization can support automated patient positioning reducing setup time, minimizing operator variability, and improving patient comfort.
 The implementation and pretrained models are publicly available at https://github.com/EytanKats/orgloc.
 keywords:
-Radiology workflow, Organs localization
+Radiology workflow automation, Automated patient table positioning, Body surface-to-3D anatomy prediction, Hybrid 2D–to-3D neural architecture.
 doi:
@@
 […]
-The growing demand for radiology examinations underscores the need to optimize scanning procedures to reduce patient wait times and improve operational efficiency.
-A critical, yet time-intensive, aspect of this process is the patient preparation and planning phase, which is essential for optimizing image quality and diagnostic accuracy (van2020cinderellas).
-This phase typically involves manual positioning of the patient followed by scout imaging to facilitate detailed geometric planning for subsequent diagnostic scans.
-In standard clinical practice, a radiologic technologist initially aligns the patient on the table, often relying on external anatomical landmarks or laser positioning guides to approximate the region of interest.
-Once the patient is positioned inside the bore, a set of low-resolution, multi-planar scout images is acquired.
-These images are then used to visually confirm the patient’s position and to manually plan the geometry of the diagnostic sequences, such as slice orientation, field-of-view, and scan range (koken2009towards).
+The growing demand for radiological examinations highlights the need to optimize scanning workflows in order to reduce patient waiting times and improve operational efficiency.
+An important yet time-consuming component of this workflow is the scan preparation and planning phase, which plays a crucial role in ensuring image quality and diagnostic reliability (van2020cinderellas).
+According to (streit2021analysis), scan preparation requires on average 5.7 minutes, corresponding to 11% of the total examination time.
+This phase typically involves manual patient positioning followed by scout imaging to enable geometric planning of the subsequent diagnostic acquisition.
+In routine clinical practice, a radiologic technologist first aligns the patient on the scanner table, often relying on external anatomical landmarks and laser positioning guides to approximate the region of interest.
+After positioning inside the scanner bore, low-resolution multi-planar scout images are acquired.
+These images are then used to verify patient positioning and to manually define the geometry of the diagnostic sequences, including slice orientation, field of view, and scan range (koken2009towards).
 Despite recent advances in radiology hardware and software, the automation of patient positioning remains a largely underexplored area compared to other aspects of scan acceleration or reconstruction (shafieizargar2023systematic; heckel2024deep).
@@
 Furthermore, variability in technologist expertise and patient anatomy can lead to inconsistent positioning, necessitating repeat localizers or even additional diagnostic scans, thereby increasing overall examination time and reducing scanner throughput (danilouchkine2005operator).
-Figure 1: Overview of the proposed automated patient positioning workflow.
-A depth sensor installed above the scanner bore captures a depth image of the patient lying on the table.
-This image is processed by a trained model that predicts the 3D bounding box of the target organ.
-The scanner table is then automatically adjusted to align the organ of interest with the center of the imaging field-of-view, reducing the need for manual repositioning and scout scans.
 Automated patient positioning systems, capable of accurately identifying and localizing the anatomical region of interest, hold considerable promise for streamlining radiology workflows by automatically adjusting the patient table to the optimal position.
 By reducing reliance on manual table adjustments and anatomical landmark estimation, these systems can minimize inter-operator variability, decrease setup time, and mitigate the risk of positioning errors, improving both imaging efficiency and diagnostic consistency across radiology departments.
-Among the emerging technological approaches in this area, RGB-D camera-based systems have shown notable potential in enhancing patient setup and positioning (booij2021automated; booij2019accuracy).
-These integrated systems use depth-sensing to generate surface maps of the patient’s body.
-By analyzing these surface contours, it becomes possible to detect and localize key anatomical landmarks, which serve as proxies for internal anatomical regions of interest.
-This information can then be used to infer the optimal imaging volume and adjust the patient table position automatically prior to scan acquisition, without the need for scout imaging.
-For example, incetan2020rgb demonstrated the utility of RGB-D sensors in automatically localizing target regions such as the head and thorax with high spatial accuracy.
-Similarly, karanam2020towards employed depth-aware landmark detection models to support robust, contactless patient positioning in MRI and CT workflows.
-In the proposed workflow, a depth sensor is integrated into the room above the scanner bore to capture a depth image of the patient lying on the table.
-This depth image is then processed by the trained model, which outputs the 3D coordinates of the bounding box enclosing the organ of interest.
-Based on this prediction, the scanner table can be automatically adjusted to ensure that the relevant anatomy is optimally centered within the imaging field-of-view, minimizing the need for manual repositioning or iterative scout imaging (Figure 1).
-This approach offers a contactless, fast, and standardized method of patient localization, which could be particularly beneficial in high-throughput clinical environments or for procedures requiring precise anatomical targeting.
-To ensure robust performance across a wide range of anatomical variations and body types, we train our model on a large-scale dataset consisting of approximately 10,000 whole-body MRI scans from the German National Cohort (NAKO) study (bamberg2022whole).
-This dataset includes subjects of diverse ages, body sizes, and physiological characteristics, enabling the model to generalize effectively across different patient populations.
-Our experimental evaluation demonstrates the effectiveness of the proposed method in accurately localizing 41 distinct internal anatomical structures, including bones and soft tissue organs, using only depth images.
-The key contributions of this work are as follows:
-  1. 1.
-    We demonstrate the feasibility and effectiveness of a learning-based framework for estimating 3D locations of internal anatomical structures solely from depth images of the patient’s body surface.
-  2. 2.
-    We introduce a convolutional neural network architecture, Pixel-to-Voxel (Pix2Vox), which maps 2D depth images to 3D volumetric segmentation masks, enabling organ localization from a single surface view.
-  3. 3.
-    We validate the generalizability of our model across a wide range of body types and anatomical variations by training and evaluating on a large-scale, population-level dataset comprising approximately 10,000 whole-body MRI scans.
-2 Related Works
-The prediction of internal anatomical structures from external body surface representations has gained increasing attention due to its potential to enable non-invasive, efficient, and patient-specific modeling.
-Early work by wu2018towards introduced a learning-based framework that generates volumetric anatomical phantoms of skeletal and pulmonary structures from triangulated meshes of the patient’s external surface.
-This study demonstrated the feasibility of inferring detailed 3D internal anatomy from external geometry using deep models trained on co-registered body surface and internal organ data.
-Intermediate 3D representations for internal anatomy estimation.
-Several approaches address the prediction of internal anatomy by first deriving an intermediate 3D representation from depth images, which is then used to infer internal structures.
+1.1 Related Works
+Surface-based patient positioning systems.
+Among the emerging technological approaches for automating patient table positioning, RGB-D camera-based systems have demonstrated significant potential for enhancing clinical workflows (booij2021automated; booij2019accuracy).
+These systems utilize depth sensing to capture the external body surface, identifying anatomical landmarks that serve as proxies for internal regions of interest.
+This information enables the localization of key structures, which can be used to infer the optimal imaging volume and adjust the patient table automatically prior to scan acquisition, potentially eliminating the need for scout imaging.
+For instance, incetan2020rgb demonstrated the utility of RGB-D sensors in localizing regions such as the head and thorax with high spatial accuracy.
+Similarly, karanam2020towards employed depth-aware landmark detection to support robust, contactless positioning in MRI and CT workflows.
+However, these approaches primarily focus on external positioning and do not explicitly infer the 3D locations or shapes of internal anatomy.
+Parametric body-model-based internal anatomy estimation.
+Several approaches address the prediction of internal anatomy by first deriving an intermediate 3D representation of the body, which is then used to infer internal structures.
 Building on the integration of statistical human body models with internal anatomy, recent research has extended the widely adopted Skinned Multi-Person Linear (SMPL) model (loper2015smpl) to incorporate internal organ representations.
@@
 (guo2022smpl) proposed SMPL-A, which augments SMPL with deformable internal organ geometries, allowing organ shapes to adapt dynamically with body pose.
-While effective in modeling pose-aware internal anatomy, SMPL-A requires a patient-specific full-body mesh to initialize internal organ shapes, limiting its generalizability and real-time applicability in clinical environments.
+While effective in modeling pose-aware internal anatomy, SMPL-A requires a patient-specific full-body mesh to initialize internal organ shapes, limiting its generalizability and applicability in clinical environments.
 Similarly, OSSO (keller2022osso) and BOSS (shetty2023boss) infer internal organ shapes and locations by optimizing anatomical plausibility within the SMPL framework based on external body surface meshes.
+However, these methods are constrained by a limited scope: OSSO focuses exclusively on the skeletal system, and BOSS models only 12 specific anatomical structures.
 Extending this paradigm, the HIT (Human Internal implicit Tissues) framework (keller2024hit) employs a deep implicit function to predict continuous volumetric tissue distributions conditioned on a parametric body surface mesh.
-HIT leverages large-scale synthetic datasets and a canonical space aligned with SMPL to infer anatomically coherent internal structures.
-Despite promising results, these SMPL-based approaches depend on accurate surface mesh reconstruction and template fitting, which are computationally intensive.
-The LOOC framework (henrich2025looc) introduces an implicit volumetric model to infer internal organ locations and approximate shapes from depth-derived surface point clouds.
-Prediction of 2D organ projections.
-Recent studies by geissler2025predicting and kats2025internal successfully utilize convolutional neural networks (CNNs) that operate directly on depth maps to predict 2D projections of organs in the coronal plane.
-These works demonstrate that CNNs can effectively extract spatial cues from external body surfaces to infer accurate organ locations.
-Prediction of 3D features from 2D depth images.
-Teixeira et al.
-(teixeira2023automated) proposed a CNN-based method that directly estimates critical scan planning parameters—such as lung centroids and patient isocenter—from raw depth images.
-This approach bypasses SMPL-based registration and operates solely on depth data, providing a lightweight and robust solution for specific imaging tasks such as isocenter calculation in thoracic MRI.
-In contrast to previous work on organ localization, our method introduces a lightweight, end-to-end framework that directly predicts explicit 3D bounding boxes and estimates shapes for 41 internal anatomical structures from a single depth image.
-Leveraging a unified Pix2Vox convolutional architecture, our approach eliminates the need for intermediate steps such as 3D surface reconstruction or canonical space registration, enabling rapid inference and practical integration into clinical workflows.
-To foster reproducibility and facilitate further research, our code and pretrained models are publicly available at https://github.com/EytanKats/orgloc.
+While HIT achieves anatomically coherent internal structures, it focuses on general tissue distributions rather than the localization of discrete, individual organs.
+Furthermore, SMPL-based approaches typically require a complete 3D surface mesh of the entire body for initialization or template registration which are not available in clinical settings.
+Estimation of internal anatomy from depth images.
+Recent studies by geissler2025predicting and kats2025internal demonstrate that convolutional neural networks operating directly on depth maps can infer three-dimensional organ locations and approximate shapes from external body surface information.
+These works highlight that depth-derived surface cues contain sufficient anatomical information to estimate internal structures without explicit geometric reconstruction.
+While these works demonstrate that external body surface information contains useful cues about internal anatomy, their predictions are limited to planar projections rather than full volumetric organ localization.
+Similarly, teixeira2023automated proposed a CNN-based method that estimates specific scan-planning parameters, such as lung centroids and patient isocenter, directly from depth images.
+Although effective for targeted task, such predictions provide only sparse anatomical descriptors and do not capture full organ extent or shape.
+A closely related concurrent study is the LOOC framework (henrich2025looc), which infers internal organ locations and approximate shapes using an implicit volumetric occupancy model conditioned on depth-derived surface point clouds.
+The approach first reconstructs a patient-specific external surface representation from depth data and subsequently predicts internal anatomy through a learned continuous occupancy field, enabling anatomically plausible volumetric estimates.
+1.2 Contribution
+This work presents a learning-based framework for estimating internal volumetric anatomical structures directly from a single anterior depth image.
+In contrast to surface-based patient positioning systems, which rely on external landmark detection or approximate region-of-interest estimation from body surface data (booij2021automated; booij2019accuracy; incetan2020rgb; karanam2020towards), our method explicitly predicts the three-dimensional locations and approximate shapes of internal anatomical structures.
+Unlike parametric body-model approaches that infer internal anatomy through template fitting or full-body surface mesh reconstruction within statistical body models (guo2022smpl; keller2022osso; shetty2023boss; keller2024hit), the proposed framework performs direct end-to-end depth-to-volume inference without intermediate body model fitting.
+Furthermore, whereas several recent depth-based approaches focus on predicting sparse anatomical descriptors such as organ centroids, isocenters, or limited projection-based organ localization (teixeira2023automated; geissler2025predicting; kats2025internal), our method produces explicit volumetric organ segmentations suitable for anatomically informed scan planning.
+Compared to the closely related concurrent LOOC framework (henrich2025looc), which employs an implicit volumetric occupancy model conditioned on reconstructed surface point clouds, our approach differs substantially in methodology by directly predicting volumetric anatomy from depth images using a hybrid 2D-to-3D convolutional neural network.
+To support reproducibility and further research, the code and pretrained models are publicly available at https://github.com/EytanKats/orgloc.
+The key contributions of this work are summarized as follows:
+  1. 1.
+    We introduce a hybrid 2D-to-3D convolutional neural network architecture (Pix2Vox) that maps a single anterior depth image to volumetric segmentations of 41 anatomical structures, providing explicit spatial information for automated radiological positioning.
+  2. 2.
+    We demonstrate large-scale training and validation on a dataset comprising 10,020 MRI-derived paired depth and volumetric anatomy samples, enabling robust modeling of anatomical variability across a diverse population.
+  3. 3.
+    We qualitatively show that a model trained purely on synthetic depth–anatomy pairs generalizes to real-world depth images of volunteers positioned in an MRI scanner, including scenarios with clothing, supporting the feasibility of practical clinical deployment.
 2 Materials and Methods
+In this study, we leverage $10,020$ whole-body MRI scans from the German National Cohort (NAKO) dataset (bamberg2022whole).
+This diverse dataset captures a broad spectrum of anatomical variability across age, sex, and body habitus, providing a foundation for training generalizable models of internal anatomy.
+As paired depth images and expert organ segmentations are not natively provided, we designed a pipeline to synthetically generate supervised training pairs directly from the MRI volumes.
+To synthesize the input data, we simulate depth sensor acquisitions by rendering the anterior contours of the patient’s body from the MRI volumes.
+This process mimics the orthographic top-down perspective of a sensor mounted above the scanner table (Section 2.2).
+To extract ground-truth 3D masks for 41 internal anatomical structures we utilize the TotalSegmentatorMRI tool (d2024totalsegmentator)(Section 2.3).
+The resulting dataset, comprising 2D depth images paired with 3D anatomical segmentations, is used to train a Pix2Vox convolutional neural network (Section 2.5).
+This architecture is optimized to map 2D surface depth information to 3D volumetric predictions of the underlying target organs.
+Once trained, the model enables a depth sensor-augmented radiology workflow, facilitating automated patient positioning through anatomical structures localization (Section 2.1).
+2.1 Radiology Workflow with Integrated Depth Sensor
+Figure 1: Overview of the proposed automated patient positioning workflow.
+A depth sensor mounted above the bore captures the patient’s surface geometry.
+Through a spatial calibration and coordinate transformation process, the depth data is mapped into the scanner’s coordinate system.
+The trained model predicts 3D organ segmentation within the scanner coordinate system.
+From this, a bounding box is derived to enable automated table translation, aligning the target anatomy with the imaging isocenter.
+In the proposed workflow, a depth sensor is mounted above the scanner bore to capture an anterior depth image of the patient on the table.
+The sensor is spatially calibrated with the MRI system, allowing the acquired depth data to be transformed into a surface point cloud within the scanner’s coordinate frame (booij2021automated; booij2019accuracy).
+This point cloud is subsequently rendered as an orthographic depth projection (Section 2.4).
+The resulting processed image is then fed into the trained model, which predicts the 3D segmentations of the anatomical structures of interest in scanner space.
+Based on these predictions, the scanner table can be automatically translated to ensure the target anatomy is optimally centered within the imaging isocenter, thereby minimizing the reliance on manual adjustments or iterative scout imaging (Figure 1).
+This approach offers a contactless, fast, and standardized method of patient localization, which could be particularly beneficial in high-throughput clinical environments.
 Figure 2: Overview of the proposed hybrid Pix2Vox network architecture.
@@
 For clarity, the batch dimension is omitted from all illustrated tensor shapes.
-In this study, we leverage 10,020 full-body MRI scans from the National Cohort Study of Germany (NAKO) dataset (bamberg2022whole).
-This large and diverse dataset enables us to capture a broad spectrum of anatomical variability across age, sex, and body habitus, providing a robust foundation for training generalizable models of internal anatomy.
-Since paired depth images and organ segmentations are not natively available in the dataset, we synthetically generate training pairs from the MRI data.
-To create depth images, we simulate depth sensor output by rendering the outer contours of the patient’s body from MRI volumes, mimicking the top-down perspective of real-world RGB-D sensors typically mounted above the scanner table.
-Simultaneously, we apply state-of-the-art automated segmentation tools to extract ground-truth 3D labels for 41 internal anatomical structures, including bones and soft tissue organs (Section 3.1).
-The resulting dataset, composed of 2D depth images paired with corresponding 3D anatomical segmentations, serves as the training input for our proposed learning pipeline.
-We devised CNN and train Pix2Vox convolutional neural network, which maps each 2D depth image to a 3D volumetric prediction of the target organ bounding boxes Section 3.2).
-3.1 Data Preparation
-To simulate depth images from MRI volumes, we begin by normalizing each 3D MRI scan to the intensity range $[0,1]$.
-Next, we apply a global intensity threshold of 0.02 to generate a binary mask that captures the patient’s body.
-Small noise components and isolated artifacts are subsequently removed using binary morphological opening with a spherical structuring element, producing a cleaner mask.
-We then extract 2D orthographic depth projections in the coronal plane, simulating the top-down view of a depth sensor positioned above the patient.
-Specifically, for each pixel in the image plane, the depth value is computed as the closest point of intersection along the anterior-posterior axis with the segmented patient’s body.
-The resulting depth maps are normalized again to the range $[0,1]$.
-To suppress non-anatomical structures such as the patient table and background noise, we apply a depth cutoff, setting all pixel values below 0.3 to 0.
-This effectively removes table artifacts and regions far below the simulated camera plane.
-Finally, to refine the depth image, we apply grayscale morphological opening, which smooths remaining artifacts while preserving significant anatomical features such as limb contours and body curvature.
-The resulting depth images offer a standardized approximation of the patient’s body surface from a fixed sensor viewpoint and serve as input to our learning framework.
-To generate high-quality anatomical reference labels, we leveraged recent state-of-the-art automated whole-body MRI segmentation tools, including TotalSegmentatorMRI (d2024totalsegmentator), MRSegmentator (hantze2024mrsegmentator), and TotalVibeSegmentator (graf2024totalvibesegmentator).
-These models, trained on large-scale MRI datasets, are capable of accurately segmenting a wide range of anatomical structures with high spatial precision.
-To further increase the number of segmented labels, the MR volumes were converted into synthetic CT volumes to utilize segmentation models available for CT data (wasserthal2023totalsegmentator; jaus2023towards), following the approach described in mensing2025mr.
-An aggregation procedure was implemented to select, for each anatomical structure, the segmentation mask with the highest accuracy.
-When a structure was available from multiple models, MR-based segmentations were prioritized over CT-based ones.
-The MRI segmentation models were ranked according to the similarity of their training data to our dataset, in the following order: TotalVibeSegmentator, MRSegmentator, and TotalSegmentatorMRI.
-For structures exclusively available from CT-based models, TotalSegmentator was preferred over Atlas DAP.
-The resulting aggregated masks encompass 137 distinct structures, including organs, vessels, bones, and tissue types.
-Given the varying quality of automated segmentations across different structures and the need for relevance to clinical workflow optimization, we performed a careful selection process.
-In consultation with domain experts, and through extensive visual inspection of the output masks, we selected 41 anatomical structures (Table 1).
-These included bones and soft-tissue organs that are both (1) of practical importance for guiding scan planning and patient positioning, and (2) reliably segmented across a broad range of patients.
-To further improve the consistency of the labels and reduce artifacts introduced by the segmentation models, we applied a series of post-processing steps.
-These included hole filling to restore anatomical completeness, and connected components analysis to eliminate spurious isolated regions or misclassified voxels.
-These refinements ensured cleaner and more robust ground-truth volumes for training the model.
+2.2 Synthetic Depth Image Generation
+To facilitate the training of a body surface-to-internal organs model, we designed a deterministic pipeline to synthesize 2D orthographic depth projections from 3D MRI volumes.
+The process transforms volumetric data $\mathcal{V}\in\mathbb{R}^{H\times W\times D}$ into depth maps $D(x,z)$
+First, MRI intensity values are normalized to the range $[0,1]$. A binary mask $M$ is generated by applying a low-intensity threshold $\epsilon=0.02$ to reduce the inherent noise:
+$M_{i,j,k}=[\mathcal{V}_{i,j,k}>0.02]$ (1)
+To remove floating artifacts and ensure structural continuity of the skin surface, we apply binary morphological opening $\gamma_{B}(M)$ using a spherical structuring element with diameter equal to 3.
+The depth image $D$ is generated via a ray-casting approach along the anterior-posterior axis.
+For each lateral and longitudinal coordinate $(x,z)$, the distance to the first encountered foreground voxel is computed:
+$d(x,z)=\text{argmax}_{y}(M_{x,y,z})$ (2)
+The raw indices are normalized by the maximum volume height $H_{max}$ and inverted so that intensity values are inversely proportional to the distance from the sensor:
+$D_{norm}(x,z)=1-\frac{d(x,z)}{H_{max}}\quad\text{for }d(x,z)>0$ (3)
+To clean signal from the scanner table and hardware artifacts, we apply a clipping threshold, zeroing all pixels with $D_{norm}<0.3$.
+High-frequency surface noise and acquisition spikes are suppressed using grayscale morphological opening with an $11\times 11$ kernel.
+To bridge the domain gap between clean synthetic depth maps and real-world clinical acquisitions, we implemented a stochastic augmentation pipeline designed to simulate the depth profile of hospital gowns, blankets, and clothing.
+The augmented depth map $D_{aug}$ is generated by modifying the synthetic depth $D$ within the patient silhouette mask $M$ as follows:
+$D_{aug}=G_{\sigma}\left(D+[B_{lines}+B_{clusters}]\cdot M\right)$ (4)
+where $G_{\sigma}$ denotes a global Gaussian smoothing kernel with $\sigma=5$, $B_{lines}$ is a line-like fold simulation and $B_{clusters}$ cluster-like noise simulation.
+Line-like Fold Simulation ($B_{lines}$).
+We simulate fabric creases by generating $N\in[5,20]$ random linear segments.
+For each segment, the orientation is sampled from $\mathcal{U}(0,2\pi)$ and the length from $\mathcal{U}(H/12,H/2)$, where $H$ is the image height.
+Each segment is assigned a randomized thickness of 141-4 pixels and blurred using a Gaussian kernel with $\sigma_{line}\in[S/20,S/5]$, where $S=100$ is a spatial scale factor.
+Folds are assigned a random polarity $p\in\{-1,1\}$ and scaled to a maximum displacement of $30$ mm.
+Cluster-like Variations ($B_{clusters}$).
+To mimic the effect of bunched blankets or hospital gowns, we generate multi-scale, low-frequency noise clusters.
+This is achieved by sampling random noise at three spatial scales $\{0.5S,S,2.0S\}$.
+For each scale, a coarse noise grid is generated and then expanded to the full image resolution using cubic interpolation.
+Each upscaled layer is refined with a Gaussian smoothing kernel where the standard deviation is scaled to the specific layer’s frequency ($\sigma=s/10$).
+This process produces a continuous, bias field that simulates bulky surface irregularities.
+The integrated cluster bias is scaled to a maximum displacement of $30$ mm.
+The augmentation process ensures that the synthetic training data exhibits the non-rigid, low-frequency spatial variations characteristic of clothes and blankets, forcing the network to prioritize global anatomical proportions over local surface textures.
+2.3 Generation of Volumetric Ground Truth Labels
+To generate anatomical reference labels, we utilized TotalSegmentatorMRI (d2024totalsegmentator), a robust deep learning framework capable of segmenting 56 distinct anatomical structures in whole-body MRI.
+To ensure the reliability of these automated segmentations as ground-truth data, we implemented a post-processing pipeline designed to improve topological consistency and eliminate segmentation artifacts.
+Specifically, we applied morphological hole filling to ensure anatomical completeness and 3D connected component analysis (CCA) to remove small isolated, spurious voxel clusters and misclassified regions.
+Due to the significant memory requirements of the 3D convolutional decoder we refined the label set to 41 target structures (Table 1).
+This selection was based on a dual-criteria approach: (1) a qualitative visual assessment of the automated segmentation accuracy across the cohort, and (2) the clinical relevance of the structure for automated scan planning and patient positioning.
+Structures with lower segmentation reliability or minimal impact on clinical workflow were excluded to optimize the model’s computational efficiency.
 Figure 3: Mean absolute Detection Offset Error (DOE) averaged across all organs for each bounding box side (left-right, anterior-posterior, superior-inferior).
-The CNN models demonstrates strong localization accuracy, particularly in the coronal (left-right) and vertical (superior-inferior) dimensions, significantly outperforming the mean model baseline.
-They achieve sub-10mm average error in the coronal plane.
-Performance degrades along the anterior-posterior axis due to limited depth cues available from the coronal input image.
-Standard deviation is depicted as error bars extending in the positive direction only.
+2.4 Processing of Real-World Depth Images
+To evaluate the robustness of our proposed model, we qualitatively assess its performance on real-world depth maps acquired from human volunteers using a Microsoft Kinect v2 sensor (Section 3.4.3).
+Raw depth frames $\mathcal{Z}\in\mathbb{R}^{U\times V}$ are first processed using a spatial Region-of-Interest (ROI) filter to isolate the subject from the scanner environment.
+This filter retains voxels within a valid depth range of $[1000,2000]$ and a lateral bounding box defined by $u\in[450,1050]$ and $v\in[400,700]$.
+Utilizing the intrinsic parameters of the depth sensor and extrinsic calibration parameters, the depth coordinates are transformed into the coordinate system of the scanner, yielding a sparse point cloud $\mathbf{P}=\{x,y,z\}$.
+The final anterior orthographic depth image is generated by raycasting $\mathbf{P}$ onto a 2D grid.
+To account for no-data pixels we employ linear interpolation to reconstruct a continuous surface.
+The resulting depth values are min-max normalized and contrast-inverted:
+$D_{real}=1.0-\frac{Z-Z_{min}}{Z_{max}-Z_{min}}$ (5)
+This transformation ensures that the region in closest proximity to the sensor is represented by high-intensity values, maintaining distribution consistency with the synthetic training data $D_{norm}$ described in Section 2.2.
 Table 1: Selected organ labels grouped by tissue type
 Soft Tissue Organs Bones
-spleen scapula left vertebrae_T12
-kidney right scapula right vertebrae_T11
-kidney left clavicula left vertebrae_T10
-liver clavicula right vertebrae_T9
-stomach femur left vertebrae_T8
-pancreas femur right vertebrae_T7
-lung right hip left vertebrae_T6
-lung left hip right vertebrae_T5
-trachea sacrum vertebrae_T4
-thyroid gland vertebrae L5 vertebrae_T3
-duodenum vertebrae L4 vertebrae_T2
-urinary bladder vertebrae L3 vertebrae_T1
+spleen scapula left vertebrae T12
+kidney right scapula right vertebrae T11
+kidney left clavicula left vertebrae T10
+liver clavicula right vertebrae T9
+stomach femur left vertebrae T8
+pancreas femur right vertebrae T7
+lung right hip left vertebrae T6
+lung left hip right vertebrae T5
+trachea sacrum vertebrae T4
+thyroid gland vertebrae L5 vertebrae T3
+duodenum vertebrae L4 vertebrae T2
+urinary bladder vertebrae L3 vertebrae T1
 aorta vertebrae L2
@@
 3 Experiments and Results
-Figure 4: Dice coefficient for volumetric organs segmentation.
-Values are averaged for bilateral organs (e.g., kidneys, lungs) and vertebral subgroups (thoracic and lumbar vertebrae).
-Pix2Vox significantly outperforms the baseline, demonstrating its ability to generate accurate, patient-specific organ segmentations from a single coronal depth image.
+Figure 4: Dice coefficient for volumetric organs segmentation. Values are averaged for bilateral organs (e.g., kidneys, lungs) and vertebral subgroups (thoracic and lumbar vertebrae).
 To train our model, we randomly selected 8,528 pairs of depth images and their corresponding segmentation masks from the preprocessed dataset ( Section 2.2).
@@
 Values are averaged for bilateral organs (e.g., kidneys, lungs) and vertebral subgroups (thoracic and lumbar vertebrae).
-Pix2Vox exhibits markedly lower surface distances compared to the mean model, indicating superior boundary accuracy and shape fidelity.
 3.2 Baselines
@@
 Figure 6: Qualitative comparison between predicted and ground truth volumetric segmentations.
-The Pix2Vox model demonstrates a strong ability to accurately infer both the shape and location of internal organs solely from a single depth image.
-The examples highlight successful predictions for both bony (vertebrae and scapula) and soft tissue organs (lung and heart) across different organ sizes and positions.
 3.3 Metrics
@@
 Bounding boxes are derived from the volumetric segmentation masks produced by the mean model, the Pix2Vox model, and the three-plane predictions of the 2.5D CNN.
-To quantify localization accuracy, we calculate the Mean Absolute Detection Offset Error (DOE), defined as the absolute distance between the corresponding faces of the predicted and ground truth bounding boxes along each axis.
+To quantify localization accuracy, we calculate the Mean Absolute Detection Offset Error (DOE), defined as the absolute distance between the corresponding faces of the predicted and ground truth bounding boxes along each axis:
+$\text{DOE}=\frac{1}{N}\sum_{i=1}^{N}|\text{face}_{i,\text{pred}}-\text{face}_{i,\text{gt}}|$ (6)
+where $\text{face}\in\{x_{min},x_{max},y_{min},y_{max},z_{min},z_{max}\}$ represents the coordinates of the six bounding box planes.
 To better understand model behavior in different spatial dimensions, DOE is reported separately for each side of the bounding box (e.g., left, right, anterior, posterior, superior, inferior).
@@
 3.4 Results
-Quantitative analysis using the Detection Offset Error (DOE) highlights the strong performance of the proposed Pix2Vox model in localizing internal organs in 3D space based solely on a single coronal depth image (Figure 3).
-Specifically, in the coronal plane, Pix2Vox demonstrates a more than twofold improvement in mean absolute DOE compared to the mean model baseline, achieving an average localization error of less than 10mm across all evaluated organs.
-However, performance degrades when localizing organs along the anterior-posterior (depth) axis, where the model shows higher variance and larger prediction errors.
-This limitation is not unexpected, given that the coronal depth image inherently provides less depth-specific information.
-Nonetheless, the ability of the CNN to infer organ depth from contextual and structural cues in a single 2D image remains noteworthy, demonstrating the model’s strong capacity for generalization.
-The 2.5D CNN approach, which integrates predictions from coronal, sagittal, and axial projections, exhibits superior localization performance in the coronal plane and achieves more robust bounding box estimation in lateral and vertical dimensions.
-However, it still underperforms in predicting organ position along the depth axis.
-This shortfall may be attributed to the limited representational capacity of the 2D projection-based architecture and the absence of a fully 3D decoding mechanism, which constrains the model’s ability to jointly reason across spatial dimensions.
-Quantitative evaluation using the Dice coefficient and ASSD)(Figure 4 and Figure 5), alongside qualitative visualizations of Pix2Vox predictions (Figure 6), reveals the model’s remarkable ability to infer internal organ shape from a single 2D depth image.
-Unlike the mean model baseline, which simply reflects anatomical averages, Pix2Vox demonstrates a substantial improvement in both overlap and surface accuracy metrics.
-This performance gap highlights the model’s capacity to move beyond population-level anatomical assumptions and instead learn spatially-aware, patient-specific volumetric representations directly from depth cues present in surface imagery.
-The model’s strong performance across diverse organ types, sizes, and locations further supports the idea that it has learned robust spatial priors, enabling generalization across a heterogeneous population.
-These findings suggest that internal organ anatomy can be successfully estimated directly from body surface depth information, highlighting its value as a non-invasive surrogate for internal imaging.
-This demonstrates the potential of integrating camera-based organ localization into radiological workflows, where such technology could enhance scan planning and optimize patient positioning.
+3.4.1 Localization Performance
+Quantitative analysis using DOE highlights the balanced 3D localization performance of the proposed Pix2Vox model across all anatomical axes, as illustrated in Figure 3.
+Within the coronal plane, Pix2Vox demonstrates a substantial improvement over the population-based mean model.
+Specifically, it achieves a mean absolute error of $7.87\pm 7.69$ mm for the right boundary and $6.87\pm 6.62$ mm for the left, significantly outperforming the baseline values of $19.86\pm 17.14$ mm and $21.23\pm 17.19$ mm, respectively.
+The 2.5D CNN approach, leveraging multi-planar projections, achieves the highest precision in lateral localization with a left-side DOE of $2.79\pm 2.93$ mm.
+While both deep learning approaches exhibit comparable performance along the superior-inferior axis, with errors ranging between $8.39$ mm and $11.36$ mm, the anterior-posterior axis remains the most challenging dimension for the 2.5D method.
+In the anterior direction, Pix2Vox yields an error of $13.29\pm 8.41$ mm, which is notably superior to the $21.31\pm 9.53$ mm error produced by the 2.5D CNN.
+Interestingly, the mean model provides a competitive baseline along the depth axis, yielding errors of $13.00\pm 10.63$ mm in the anterior direction and $12.90\pm 12.48$ mm in the posterior direction.
+Figure 7: Qualitative evaluation on real-world depth images of volunteers positioned in an MRI scanner, following the setup described in Section 2.
+The predicted internal structures are displayed as volumetric segmentations projected onto the coronal plane.
+From left to right: (1) input depth image; (2) overlay of lungs, kidneys, and urinary bladder; (3) overlay of trachea, heart, spleen, liver, and pancreas; (4) overlay of thyroid gland, aorta, stomach, and duodenum; and (5) volumetric segmentation of the skeletal system, shown in a different perspective for each subject.
+3.4.2 Segmentation and Shape Fidelity
+The volumetric reconstruction performance was further evaluated using the Dice similarity coefficient and ASSD.
+The Pix2Vox model achieved an overall mean Dice score of $0.48\pm 0.20$, representing a clear advancement over the population-based mean model baseline of $0.34\pm 0.21$.
+Anatomical analysis reveals that large organs with a high correlation to the body surface exhibit the most robust reconstruction.
+Specifically, the lungs and liver achieved Dice scores of $0.78\pm 0.04$ and $0.71\pm 0.07$, respectively.
+In contrast, smaller or highly variable structures, such as the thyroid gland at $0.24\pm 0.15$ and the trachea at $0.32\pm 0.19$, presented greater challenges for surface-based volumetric estimation.
+Skeletal structures also showed competitive performance, with the femur at $0.62\pm 0.10$ and the sacrum at $0.61\pm 0.15$ yielding stable overlap metrics.
+The surface fidelity metrics follow a similar trend.
+Pix2Vox reduced the overall mean ASSD to $8.34\pm 8.11$ mm, compared to $13.17\pm 10.84$ mm for the mean model.
+The highest surface accuracy was observed for the scapula and the hip, with errors of $5.96\pm 1.65$ mm and $6.08\pm 1.27$ mm, respectively.
+Notably, while some internal organs exhibit lower Dice scores due to their small volume, the corresponding ASSD values remain relatively low, as seen with the $6.58$ mm error for the thyroid gland.
+3.4.3 Evaluation on Real-World Depth Images
+To assess the translational potential of our framework, we evaluated the trained model on depth maps acquired from volunteers using a Microsoft Kinect v2 sensor.
+As illustrated in Figure 7, the model reconstructs internal anatomical structures that appear spatially consistent with the subject’s external surface manifold.
+The predicted organ locations and global orientations align with expected human anatomy, suggesting that the model is robust to the real-world depth images.
+However, detailed visual inspection reveals specific reconstruction artifacts inherent to surface-only volumetric estimation.
+In the skeletal overlays, particularly in the middle image of Figure 7, morphological artifacts are noticeable in the femoral regions, appearing as unusually narrowing towards bottom of the image.
+Furthermore, the lung segmentations in the middle and bottom rows exhibit localized peak-like artifacts.
+While the overall abdominal arrangement is preserved, the morphology of the liver in middle image shows a loss of characteristic lobar sharpness.
 4 Discussion
-In this study, we introduce an approach for automated patient positioning that leverages body surface depth images to estimate the 3D position and shape of internal organs.
-Our quantitative and qualitative results demonstrate the feasibility of using a single depth image to approximate the internal anatomical layout.
-The proposed Pix2Vox model successfully learns to extract cues from surface morphology, such as torso shape, curvature, and contour depth, to predict the volumetric extent and spatial orientation of internal organs.
-The model’s strong localization accuracy across multiple organs a promising direction to support scan planning and optimize table positioning without the need for external markers or scout imaging.
-Automated positioning could reduce reliance on trial-and-error adjustments by radiology technicians, potentially decreasing scan time, improving reproducibility, and reducing inter-operator variability.
-Additionally, the model’s capacity to generalize across diverse patient anatomies, as evidenced by its performance on a large and heterogeneous test set, underlines its robustness in dealing with physiological variation.
-Despite these encouraging findings, several limitations and opportunities for future research must be acknowledged:
-  • Domain gap between synthetic and real-world depth data: The depth images used in this study were synthetically generated from MRI-derived segmentations and surface reconstructions, which may not fully capture the variability present in real-world depth camera outputs.
-In clinical environments, additional sources of noise, such as occlusions, limited spatial resolution, lighting artifacts, and sensor-specific distortions, can affect image quality.
-Furthermore, patients may be covered with blankets, wear hospital gowns, or use assistive devices, all of which can obscure key body surface features.
+The quantitative and qualitative results indicate that 3D internal anatomy can be estimated from 2D surface depth information, though the degree of accuracy is contingent upon the specific organ of interest and the anatomical axis.
+Analysis of the Detection Offset Error reveals distinct trade-offs between the evaluated architectures.
+While the Pix2Vox model provided more balanced 3D localization across all axes, the 2.5D CNN achieved the highest precision in the lateral left-right dimensions.
+This is likely attributable to its multi-planar projection approach, which excels at identifying boundary edges in 2D space.
+However, the performance degradation of the 2.5D CNN along the anterior-posterior axis suggests that the absence of a dedicated 3D decoding mechanism limits its ability to resolve depth-specific structural cues.
+The lower performance in the anterior-posterior direction observed for both deep learning approaches is not unexpected, as a single-view coronal depth map provides limited information regarding the posterior extent of internal organs.
+The disparity in Dice scores between large organs, such as the lungs, and smaller structures, such as the thyroid gland, suggests that volumetric overlap is highly dependent on the absolute size of the target.
+However, the relatively low ASSD for small organs indicates that the model accurately localizes the anatomical center of these structures even when it cannot precisely resolve their morphological boundaries.
+4.1 Real-World Generalization and Artifacts
+The evaluation on real-world Kinect v2 data (Section 3.4.3) demonstrates that the model is robust to real-world depth acquisition, but it also highlights inherent limitations in surface-to-volume inference.
+The narrowing artifacts observed in the femoral regions and the peak-like artifacts in the lung segmentations suggest a sensitivity to surface curvature and sensor noise.
+The loss of lobar sharpness in the liver morphology further indicates that internal soft-tissue boundaries are not fully constrained by the surface manifold.
+These discrepancies suggest that while global positioning is effectively captured, fine-grained boundary reconstruction of non-surface-correlated tissues remains a challenge.
+4.2 Limitations and Clinical Considerations
+Several technical constraints must be addressed for clinical implementation:
+  • Domain Gap: The depth images used in this study were synthetically generated from MRI-derived segmentations and surface reconstructions, which may not fully capture the variability present in real-world depth camera outputs.
+In clinical environments, additional sources of noise, such as occlusions, limited spatial resolution, and sensor-specific distortions, can affect image quality.
+Furthermore, patients may be covered with thick blankets or use assistive devices that obscure key body surface features.
 Bridging this domain gap is crucial for real-world applicability and may require domain adaptation techniques or model fine-tuning using real sensor-acquired data.
-  • Assumption of consistent patient pose and orientation: The training data used in this work consisted of patients in a standardized supine position with minimal variation in posture.
-Although position-based data augmentations were employed during training to increase robustness, real-world scenarios often involve a wider range of patient poses due to comfort needs or mobility restrictions.
-The current model may be sensitive to these variations, and further research is needed to evaluate and improve its generalizability to diverse body orientations and clinical settings.
+  • Positional Constraints: The training data used in this work consisted of patients in a standardized supine position with minimal variation in posture.
+Although geometric data augmentations were employed during training to increase robustness, real-world scenarios often involve a wider range of patient poses due to comfort needs or mobility restrictions.
+The current model may be sensitive to these variations, and further research is needed to improve its generalizability to diverse body orientations.
+In conclusion, the results demonstrate the technical feasibility of using depth-based volumetric estimation for automated isocenter alignment.
+While the precision is promising for the automated table positioning use-case, the observed morphological artifacts suggest that surface data alone may be insufficient for high-fidelity reconstruction of all anatomical structures.
 […]
@@
 […]
-    The work follows appropriate ethical standards in conducting research and writing the manuscript, following all applicable laws and regulations regarding treatment of animals or human subjects.
+The work follows appropriate ethical standards in conducting research and writing the manuscript, following all applicable laws and regulations regarding treatment of human subjects.
 […]
@@
 Access is subject to approval by the NAKO Data Use and Access Committee and compliance with data protection regulations.
+Appendix A Detailed Performance Metrics
+A detailed breakdown of performance metrics for each organ. Figures 1013 present the evaluation across all 41 anatomical structures.
+Figure 8: Dice similarity coefficient for internal organs.
+Figure 9: Dice similarity coefficient for internal organs.
+Figure 10: Surface distance metrics for internal organs.
+Figure 11: Surface distance metrics for internal organs.
+Figure 12: Detection offset for internal organs.
+Figure 13: Detection offset for internal organs.